1000 raffle tickets are sold at $3 each. One grand prize of $100 and two consolation prizes of $75 will be awarded. Find Jake’s expectation if he purchases one ticket.
A) -$2.75 B) -$3.00 C) $0.24 D) $2.75 E) $3.00

Answer by Edwin McCravy(20086)

1000 raffle tickets are sold at $3 each. One grand prize of $100 and two consolation prizes of $75 will be awarded. Find Jake’s expectation if he purchases one ticket.
A) -$2.75 B) -$3.00 C) $0.24 D) $2.75 E) $3.00


There are three possibilities:

1. Jake wins the grand prize of $100, but since he
   paid $3, that amounts to winning only +$97.
2. Jake wins one of the two consolation prizes of $75,
   but since he paid $3, that amounts to winning +$72.
3. Jake wins 0$, and since he paid $3, that amounts to 
   losing $3, which can be thought as "winning" -$3. 
  (Losing is considered "winning" a negative number 
   of dollars.)

The probability that 1 will occur is   1/1000 or 0.001
The probability that 2 will occur is   2/1000 or 0.002
The probability that 3 will occur is 997/1000 or 0.997
  
Winnings:   Probability     Winnings × Probability
    X           P                  X × P 
1. +$97   |   0.001      |       $97 × 0.001 = +$0.097
2. +$72   |   0.002      |       $72 × 0.002 = +$0.144
3.  -$3   !   0.997      |       -$3 × 0.997 = -$2.991
-----------------------------------------------------                  
  Totals:     1.000                            -$2.75   
   

The concept is this: If this same raffle were held 
many times, then if Jake bought one ticket every time,
and counted up all his winnings and losses, he would 
average losing $2.75 per game, and this is considered 
to be the negative number -$2.75.   

Edwin